# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ H.2 News Generation
Step 3: Human post-editing and quality assurance
The AI-generated draft is reviewed and refined by human annotators to enhance readability, ensure alignment with journalistic standards, and eliminate any hallucinated or unverifiable claims. This step also ensures the final article is copyright-compliant and fully owned by the authors.
H.3 Question Design
@@
H.4 Annotation Guideline
H.4.1 Task Description
@@ H.4.2 Annotation Rules
H.5 Annotator Demography
The construction of our multilingual financial benchmark relies on the deep domain expertise of a team of highly qualified annotators with strong backgrounds in finance, economics, and data science. Their interdisciplinary training and multilingual fluency ensure accurate and contextually grounded annotations across both financial statements and cross-lingual news articles. This is especially critical for our dataset, which spans multiple filings across companies, each annotated with eight analytically demanding questions that require close reading, financial reasoning, and linguistic nuance.
One annotator, currently working as a senior analyst at a major U.S. financial institution, holds a master’s degree in Business Analytics from a leading Ivy League university and a bachelor’s degree in Statistics and Economics. Their background includes research on large language models (LLMs), financial data analysis, and economics, enabling precise annotation of complex financial information. Fluent in Chinese and experienced in multilingual reasoning, she brings a high level of rigor to aligning financial data with multilingual news content.
@@ H.5 Annotator Demography
The team is further strengthened by an annotator currently serving as a Research Associate at the Asian Development Bank Institute (ADBI) in Tokyo. They hold a Ph.D. in Financial Economics from a leading Australian university, where their research focused on mergers and acquisitions, machine learning, and firm behavior. They also earned an M.Phil. in Finance and a Master’s in Accounting and Finance from the same institution. Their experience includes academic positions at universities in both Australia and China, along with multiple publications and conference presentations at venues such as the Pacific-Basin Finance Journal and the 2023 Chicago FMA Conference. They have additionally contributed to the field as an editorial assistant, ad hoc referee, and research assistant, bringing a deep academic foundation to the annotation process.
@@ H.7 Validation Guideline
H.8 Annotation Agreement Metric
To account for the annotation agreement in our consistency rating task, we report Gwet’s AC1 [18], a chance-corrected agreement coefficient that is more robust to prevalence and marginal distribution issues than Cohen’s Kappa. The AC1 statistic is defined as:
[…]
@@ H.8 Annotation Agreement Metric
H.9 Instruction Data Conversion
@@
Appendix I Model Openness Framework (MOF)
@@
To systematically assess the openness and completeness of the models listed above, we adopt the Model Openness Framework (MOF) [75], a three-tier ranked classification system designed for machine learning models. The MOF defines 17 components spanning the model development lifecycle and categorizes models into three hierarchical classes: “Class III - Open Model”, “Class II - Open Tooling”, and “Class I - Open Science”, with each level subsuming the requirements of the preceding one. Class III represents the minimal threshold of openness. Although many models are advertised as “open”, few satisfy even the Class III criteria. Most release only partial artifacts, such as model architecture, weights, model cards, or technical reports, without permissive open-source licenses. Consequently, downstream use, modification, and redistribution are often legally constrained, posing potential legal and compliance risks. Among the models in our benchmark, most fall under Class III, with only llava-v1.6-vicuna-13b qualifying for Class II.
Appendix J Open Multilingual and Multimodal Financial LLM Leaderboard
@@
To promote transparency, openness, and community engagement, we have developed an interactive leaderboard for MultiFinBen [^8] https://huggingface.co/spaces/TheFinAI/Open-FinLLM-Leaderboard (Figure5). The leaderboard displays model performance across all benchmark tasks and includes Model Openness Framework (MOF) [75] tags, which provide structured metadata about each model’s openness, accessibility, and reproducibility. Users can filter, compare, and explore submissions based on performance metrics and openness levels. The leaderboard is continuously updated and publicly accessible to encourage standardized and responsible evaluation practices in financial AI.
Figure 5: The MultiFinBen leaderboard.
Appendix K Author Contribution
The author contributions are summarized below:
  • Science Leadership: Kaleb E Smith, Meikang Qiu, Arman Cohan, Xiao-Yang Liu, Jimin Huang, Alejandro Lopez-Lira, Xi Chen, Junichi Tsujii, Jian-Yun Nie, Sophia Ananiadou, Qianqian Xie
  • Contributors: Xueqing Peng, Lingfei Qian, Yan Wang, Ruoyu Xiang, Yueru He, Yang Ren, Mingyang Jiang, Jeff Zhao, Huan He, Yi Han, Yun Feng, Yuechen Jiang, Yupeng Cao, Haohang Li, Yangyang Yu, Xiaoyu Wang, Penglei Gao, Shengyuan Lin, Keyi Wang, Shanshan Yang, Yilun Zhao, Zhiwei Liu, Peng Lu, Jerry Huang, Suyuchen Wang, Triantafillos Papadopoulos, Polydoros Giannouris, Efstathia Soufleri, Nuo Chen, Guojun Xiong, Zhiyang Deng, Yijia Zhao, Mingquan Lin

### v3

@@ H.3 News Generation
Step 3: Human post-editing and quality assurance.
The AI-generated draft is subsequently reviewed by human annotators to improve readability, ensure adherence to journalistic standards, and remove any hallucinated or unverifiable content. This step guarantees that the final article is copyright-compliant and fully authored by the benchmark creators.
H.4 Question Design
@@
H.5 Reasoning Trajectory
PolyFiQA-Easy:
  • Q: What trends can be observed in the company’s revenue amount over the past few years? Please quote the piece(s) of news aligned with the finding(s), if any.
    Steps:
    1. 1.
      Extract revenue amounts over the past few years.
  • Q: How does the company’s balance sheet reflect its financial health in terms of total amount of current assets and the ratio of total liability divided by total equity? Please provide only the final result(s); do not include any calculation steps. Please quote the piece(s) of news aligned with the finding(s), if any.
    Steps:
    1. 1.
      Extract current assets, total liability, and total equity amounts.
    2. 2.
      Calculate the ratio (total liability divided by total equity).
  • Q: Are there any significant changes or irregularities in the amount of operating, investing and financing cash flow? Please quote the piece(s) of news aligned with the finding(s), if any.
    Steps:
    1. 1.
      Extract all three cash flow components (operating, investing and financing cash flow amounts).
  • Q: What is the company’s R&D ratio (R&D divided by revenue)? Please provide only the final result(s); do not include any calculation steps. Please quote the piece(s) of news aligned with the finding(s), if any.
    Steps:
    1. 1.
      Extract R&D and revenue amounts.
    2. 2.
      Calculate R&D ratio (R&D amount divided by revenue amount).
PolyFiQA-Expert:
  • Q: Please list the top three focuses on revenue from news. Please quote all the relevant financial information from the financial statements supporting the finding(s), if any.
    Steps:
    1. 1.
      Extract revenue for all focuses.
    2. 2.
      Rank the revenue amounts of different focuses from high to low.
    3. 3.
      Select the top three.
  • Q: How is the company allocating capital (e.g., investments, share repurchases, dividends) based on its performance and market outlook from the news? Please quote all the relevant financial information from the financial statements supporting the finding(s), if any.
    Steps:
    1. 1.
      Extract capital allocation items (investments, dividends, etc).
    2. 2.
      Extract market outlook information from news.
    3. 3.
      Analyze capital amounts and market outlook information together.
  • Q: What is the company’s strategy on maintaining profit margins from the news? Please quote all the relevant financial information from the financial statements supporting the finding(s), if any.
    Steps:
    1. 1.
      Extract net profit and revenue.
    2. 2.
      Calculate profit margin (net profit divided by revenue amount).
    3. 3.
      Extract strategy information.
    4. 4.
      Analyze profit margin ratio and strategy information.
  • Q: What are the company’s capital expenditures and their strategic significance from the news? Please quote all the relevant financial information from the financial statements supporting the finding(s), if any.
    Steps:
    1. 1.
      Extract capital expenditures.
    2. 2.
      Extract strategic context.
    3. 3.
      Analyze capital expenditures amounts and strategy information together.
H.6 Annotation Guideline
H.6.1 Task Description
@@ H.6.2 Annotation Rules
H.7 Annotator Demography
The construction of our 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: relies on the deep domain expertise of a team of highly qualified annotators with strong backgrounds in finance, economics, and data science. Their interdisciplinary training and multilingual fluency ensure accurate and contextually grounded annotations across both financial statements and cross-lingual news articles.
One annotator, currently working as a senior analyst at a major U.S. financial institution, holds a master’s degree in Business Analytics from a leading Ivy League university and a bachelor’s degree in Statistics and Economics. Their background includes research on LLMs, financial data analysis, and economics, enabling precise annotation of complex financial information. Fluent in Chinese and experienced in multilingual reasoning, they bring a high level of rigor to aligning financial data with multilingual news content.
@@ H.7 Annotator Demography
The team is further strengthened by an annotator currently serving as a Research Associate at a major international financial research institute in Tokyo. They hold a Ph.D. in Financial Economics from a leading Australian university, with research spanning mergers and acquisitions, machine learning, and firm behavior. Their background includes prior academic positions in both Australia and China, multiple peer-reviewed publications in finance journals, and presentations at major international conferences. They also have experience as an editorial assistant, ad hoc referee, and research assistant, contributing a strong academic foundation to the annotation process.
@@ H.9 Validation Guideline
H.10 Annotation Agreement Metric
To account for the annotation agreement in our consistency rating task, we report Gwet’s AC1 Gwet (2014), a chance-corrected agreement coefficient that is more robust to prevalence and marginal distribution issues than Cohen’s Kappa. The AC1 statistic is defined as:
$\mathrm{AC1}=\frac{P_{o}-P_{e}}{1-P_{e}},$ (2)
@@ H.10 Annotation Agreement Metric
H.11 Instruction Data Conversion
@@
Appendix I Error Analysis
I.1 PolyFiQA-Easy and PolyFiQA-Expert
Expert reviewers analyzed model outputs from GPT-4o and LLaMA3.1-70B-Instruct, identifying information extraction failures as the most common error source. Models often cited incorrect financial fields (e.g., using total assets instead of cash flow, or retained earnings instead of net profit margin), reflecting confusion over semantically similar numeric entries. In some cases, models returned “none” when they failed to locate relevant information.
These errors indicate persistent challenges in retrieving and interpreting structured financial data, especially in multilingual and cross-document settings. The significance of this task lies in its real-world complexity: financial professionals routinely integrate structured filings and multilingual news to form investment insights. PolyFiQA-Easy and PolyFiQA-Expert, as the first benchmark to capture this multilingual, cross-document reasoning challenge, provides a high-fidelity testbed for evaluating and advancing LLMs toward realistic financial QA capabilities.
I.2 EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR
Expert review of GPT-4o’s outputs identified five major error types:
  1. 1.
    Omission of tables and charts, especially in complex or multi-column layouts, leading to loss of key financial content;
  2. 2.
    Hallucinations, where models generated plausible but ungrounded content;
  3. 3.
    Numeric misinterpretation, including rounding errors, dropped digits, and misreading of formatting conventions;
  4. 4.
    Edge-based omissions, where headers, footers, or marginal content was ignored, possibly due to incomplete OCR bounding or attention bias;
  5. 5.
    Skipping bracketed or vertical text, which often contained essential qualifiers or context.
These diverse and persistent error types underscore the real-world difficulty and importance of our OCR benchmark. Accurate, layout-aware extraction of financial content is essential for downstream tasks such as auditing, reporting, and compliance. With four novel OCR datasets (EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR), our benchmark supports fine-grained diagnosis of model limitations and advances the frontier of multimodal financial NLP.
Appendix J Human Evaluation of OCR Outputs
Automatic metrics such as ROUGE-1 may not fully capture factual consistency in long-form OCR outputs. To complement quantitative evaluation, we conducted a manual factuality assessment of GPT-4o’s OCR results following a structured annotation guideline. Inter-annotator agreement (IAA) was computed to ensure reliability.
J.1 Factuality Scoring Guideline
  • 1 (Bad): Multiple factual inaccuracies, such as incorrect company names, locations, or numerical data.
  • 2 (Okay): Generally accurate but contains omissions or minor inconsistencies in the extracted content.
  • 3 (Good): Mostly accurate with small formatting deviations, such as missed page numbers.
  • 4 (Excellent): Fully faithful to the source document, with all content and structure accurately preserved.
J.2 Human Evaluation
Dataset Mean Factuality Score Cohen’s $\kappa$ Krippendorff’s $\alpha$
EnglishOCR (n=256) 2.224 0.752 0.841
SpanishOCR (n=50) 2.865 0.886 0.930
GreekOCR (n=100) 2.865 0.886 0.930
Table 8: Human factuality evaluation of OCR outputs from GPT-4o.
This evaluation highlights the limitations of automatic similarity-based metrics and demonstrates the need for manual factual assessment in layout-rich, domain-specific OCR tasks.
Appendix K Model Openness Framework (MOF)
@@ I.2 EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR
To systematically assess the openness and completeness of the models listed above, we adopt the Model Openness Framework (MOF) White et al. (2024b), a three-tier ranked classification system designed for machine learning models. The MOF defines 17 components spanning the model development lifecycle and categorizes models into three hierarchical classes: “Class III - Open Model”, “Class II - Open Tooling”, and “Class I - Open Science”, with each level subsuming the requirements of the preceding one. Class III represents the minimal threshold of openness. Although many models are advertised as “open”, few satisfy even the Class III criteria. Most release only partial artifacts, such as model architecture, weights, model cards, or technical reports, without permissive open-source licenses. Consequently, downstream use, modification, and redistribution are often legally constrained, posing potential legal and compliance risks. Among the models in our benchmark, most fall under Class III, with only llava-v1.6-vicuna-13b qualifying for Class II.
Appendix L Open Multilingual and Multimodal Financial LLM Leaderboard
@@
To promote transparency, openness, and community engagement, we have developed an interactive leaderboard for 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: (Figure 10). The leaderboard displays model performance across all benchmark tasks and includes Model Openness Framework (MOF) White et al. (2024b) tags, which provide structured metadata about each model’s openness, accessibility, and reproducibility. Users can filter, compare, and explore submissions based on performance metrics and openness levels. The leaderboard is continuously updated and publicly accessible to encourage standardized and responsible evaluation practices in financial AI.
Figure 10: The 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: leaderboard.